is there a solution for y=x^+2 x+7
Answer by rothauserc(4718)

y = x^2 +2x +7
:
a solution is
:
x^2 +2x +7 = 0
:
complete the square
:
x^2 +2x +1 = -7 +1
:
(x+1)^2 = -6
:
x+1 = square root(-6)
:
*************************
x = -1 +i*square root(6)
:
x = -1 -i*square root(6)
:
y = x^2 +2x +7 does not cross the x axis so it has no zeros, the solutions for x are complex
************************
:
note that i = square root(-1)
: